The radius of sphere 5 cm density 10500 kg/m3 d1=20 cm d2= 40 cm. Calculate the mass of the sphere ?

Grade:10

2 Answers

Rakshit Acharjee
6 years ago
A water vessel of mass 300kg and volume 900 m3 floats ofn water and while floating a buoyant force of 2.94*10^6 N acts on it. Keeping the volume same, the vessel is tried to use as submerged submarine. What part of the volume of the vessel remains submerged when the vessel was floating on water ?
Anand Mishar
6 years ago
radius (r)=5cm
1m=100cm
1cm=1/100cm
r in meter 5/100 =1/20 m 
v=(4/3)*pi*(r*r*r)
formula of volume for sphaire v=(4/3)*(22/7)*(1/20)*(1/20)*(1/20)
density d=m/v
m=d*v
d=10500 kg/m3
m=2.004545